OpenStudy (anonymous):

What is (-3) raised to the to the power of 3

OpenStudy (anonymous):

It is \[-3 \times -3 \times -3\]

OpenStudy (anonymous):

or (-3)(-3)(-3) going similar to polpak and yes it was -27

OpenStudy (anonymous):

\[^{(-5)^{3}}\]

OpenStudy (anonymous):

\[-5\times -5 \times -5\]